eClipSe:

The value of #((x-y)/z+(y-z)/x+(z-x)/y)(z/(x-y)+x/(y-z)+y/(z-x))# is ? when #x+y+z=0#

ratnakermehta20:

The Numerator (NR.) of the Expression (Exp.) is, The NR.=(x-y)/z+(y-z)/x+(z-x)/y, ={xy(x-y)+yz(y-z)+zx(z-x)} /(xyz), ={x2y-xy2+y2z-yz2+z2x-zx2}/(xyz), ={x2y-zx2-xy2+xz2+y2z-yz2}/(xyz), ={x2(y-z)-x(y2-z2)+yz(y-z)}/(xyz), ={ x2(y-z)-x(y-z)(y+z) +yz(y-z)}/(xyz), =[(y-z){x2-x(y+z)+yz}]/(xyz), =[(y-z){x2-xy-xz+yz}]/(xyz), =[(y-z){x(x-y)-z(x-y)}]/(xyz), Hence, the NR.={-(y-z)(x-y)(z-x)}/(xyz). Similarly, the Denominator (DR.) of the Exp. can be shown, DR=-{z(z-x)(z-y)+y(y-x)(y-z)+x(x-y)(x-z)}/{(y-z)(x-y)(z-x)}, =-{z(z2-(x+y)z+xy)+y(y2-(z+x)y+xz)+x(x2-(y+z)x+yz)}/{(y-z)(x-y)(z-x)}. Here, from x+y+z=0, we get, x+y=-z, y+z=-x, z+x=-y. Hence, z2-(x+y)z+xy=z2+z2+xy=2z2+xy, y2-(y+z)x+zx=2y2+zx, x2-(y+z)x+yz=2x2+yz. Therefore, DR.={-z(2z2+xy)-y(2y2+zx)-z(2x2+yz)}/{(y-z)(x-y)(z-x)}, =-{2z3+2y3+2x3+3xyz}/{(y-z)(x-y)(z-x)}. Now, recall that, when x+y+z=0, x3+y3+z3=3xyz. Hence, DR.=-{2(3xyz)+3xyz}/{(y-z)(x-y)(z-x)}. Thus, DR.=-(9xyz)/{(y-z)(x-y)(z-x)}. Therefore, the required value=(NR.)/(DR.)=9.

AdiGirl:

woah

jhonyy9:

@ratnakermehta20 very nice work !!!